OpenStudy (anonymous):

x5 - 16x3 + 8x2 - 128 x3(x2 - 16) + 8(x2 - 16) (x2 -16)(x3 + 8)

OpenStudy (anonymous):

Has the polynomial below been factored correctly and completely? Explain your answer.

OpenStudy (anonymous):

Yes, it had been factored correctly. One easy way to check this would be to multiply out the last set of terms, and see if such equals the very first expression.

OpenStudy (anonymous):

Okay thank you very much

OpenStudy (anonymous):

Oh, wait, but it has not been factored completely! Sorry, didn't catch that last part...

OpenStudy (anonymous):

oh whats wrong with it?

OpenStudy (anonymous):

Because, by the difference of squares we know that: \[ a^2-b^2=(a+b)(a-b) \]Try finding out which part is still missing

OpenStudy (anonymous):

tbh i have no idea. thats why i'm on here i don't get it

OpenStudy (anonymous):

Well, all right, so, we have the last term equals: \[ (x^2-16)(x^3+8) \]We know that: \[ a^2-b^2=(a-b)(a+b) \]And that \(x^2\) is a square, while \(16\) is also a square, thus: \[ x^2-16=(x+4)(x-4) \]We also notice that, the last term is: \[ x^3+8=x^3+2^3 \]So, both are cubes. If we use: \[ a^3+b^3=(a+b)(a^2-ab+b^2) \]We then get: \[ x^3+2^3=(x+2)(x^2-2x+4) \]So, the expression factors down *completely* AND *correctly* to: \[ (x+8)(x-8)(x+2)(x^2-2x+4) \]

OpenStudy (anonymous):

ohhh okay thanks a lot for helping me

OpenStudy (anonymous):

And, I messed up.. again, ugh, it's \((x-4)(x+4)(x+2)(x^2-2x+4)\)

OpenStudy (anonymous):

i figured that out tho

OpenStudy (anonymous):

Oh, then, all right!